The points A(3,4), B(-6,-2) and C(5,1)are the vertices of a triangle on a graphic designers template.
Find the equation of the medin of triangle ABC from A(3,4)

Answer by venugopalramana(3286)

he points A(3,4), B(-6,-2) and C(5,1)are the vertices of a triangle on a graphic designers template.
Find the equation of the medin of triangle ABC from A(3,4)
LET D BE MID POINT OF BC
HENCE D IS (-6+5)/2=-1/2 AND (-2+1)/2=-1/2
EQN OF AD THE MEDIAN IS
(Y-4)/(4+0.5)=(X-3)/(3+0.5)
3.5(Y-4)=4.5(X-3)
7(Y-4)=9(X-3)
7Y-28=9X-27
7Y-9X-1=0